Mary put in a total of 16 1/2 hours mowing lawns during 5 days of the past week. How long was her average work day?
Answer by stanbon(75887)

Mary put in a total of 16 1/2 hours mowing lawns during 5 days of the past week. How long was her average work day?
-------
(16 1/2)/5 = (33/2)/5 = 33/10 = 3 1/3 hrs per day
==================
Cheers,
Stan H.